# Large Language Models for Constructing and Optimizing Machine Learning Workflows: A Survey

### v1

@@ 1 Introduction
These limitations highlight the need for more advanced solutions, where LLMs demonstrate significant potential by incorporating prior knowledge, facilitating human-AI collaboration, and producing more interpretable outcomes. For instance, in hyperparameter optimization Zhang et al. (2023c); Liu et al. (2024a), LLMs can leverage historical data and domain-specific insights to predict optimal configurations, thereby enhancing model performance and reducing the reliance on exhaustive trial-and-error methods. Furthermore, the sophisticated NLP capabilities of LLMs enable them to act as interactive agents or chatbots, capable of generating and adjusting AutoML code based on contextual information provided by users Dakhel et al. (2023); Arteaga Garcia et al. (2024). Crucially, in the process of conversational construction of ML workflows—such as feature engineering Hollmann et al. (2024)—LLMs have shown the ability to generate human-readable and explainable features, offering a level of transparency and guidance that surpasses many traditional black-box AutoML systems Zhang et al. (2024a); Nam et al. (2024).
However, integrating LLMs into ML workflows also presents several challenges. Issues such as reasoning hallucinations, ethical concerns, and the substantial computational demands of deploying large-scale models remain significant barriers to their widespread adoption Bommasani et al. (2021); Hollmann et al. (2024); Yao et al. (2024a). Additionally, while LLMs have demonstrated impressive results in specific tasks, their effectiveness across the entire spectrum of ML workflow stages still requires thorough investigation Zhang et al. (2023b).
This survey provides a comprehensive overview of the current state of research on the application of LLMs in constructing and optimizing ML workflows. To the best of our knowledge, it is the first survey to systematically address every stage of the ML workflow (Fig. 1), distinguishing it from previous reviews, such as Tornede et al. (2023a), which primarily explore the broader opportunities of an integration of LLMs and AutoML. We organize recent papers based on the specific steps in which LLMs have been utilized, as summarized in Table 1. Notably, our focus is on capturing the breadth and possibility of research within the specific context of constructing and optimizing ML workflows, rather than providing an exhaustive examination of all LLM-related methods at each individual stage.
@@ 3.1.1 Data Acquisition
In the context of dataset recommendation, some systems harness LLMs to suggest datasets that align with specific task needs. AutoMMLab Yang et al. (2024) employs LLMs for natural language interaction with users, utilizing a "dataset zoo" to recommend or retrieve relevant datasets based on task descriptions. This automated selection helps tailor custom datasets to the unique requirements of specialized machine learning tasks. LLMs are also used for dataset summarization, aiding users in understanding data structure and content before further processing. The Virtual Interactive Data Scientist (VIDS) framework Hassan et al. (2023), for example, features ChatGPT-powered agents that extract detailed insights, including dataset structures, column details, and visualization suggestions. Similarly, JarviX Liu et al. (2023) incorporates an "Insight" component to automatically gather structured data information—such as column names, types, and statistics—and generate LLM-driven summary reports, giving users a quick overview of key dataset characteristics.
@@ 3.1.3 Data Augmentation
Data augmentation refers to techniques used to artificially increase the size of the training dataset by creating modified versions of existing data. This step is particularly useful for enhancing model robustness, improving generalization, and avoiding overfitting, especially in scenarios where data is scarce.
LLMs contribute to data augmentation by generating synthetic data, enriching existing datasets, or suggesting transformations that improve model generalization. Language models, when trained on pairs of data (e.g., images) and corresponding descriptive texts (e.g., captions), can learn the association between the textual descriptions and the underlying data. The trained model can then be used to generate desired images based on descriptive text inputs Mumuni and Mumuni (2024). Recent research highlights the data generation capabilities of Pre-trained Language Models (PLMs) to produce synthetic data for training target models Gao et al. (2023); Yu et al. (2024). To mitigate the inherent data distribution bias in synthetic datasets generated by a single PLM, FuseGen Zou et al. (2024) collaboratively leverages multiple PLMs to generate higher-quality synthetic datasets without incurring additional queries to individual PLMs, enhancing the reliability of synthetic data. Recently, NVIDIA’s Nemotron-4 340B model family Adler et al. (2024) has been incorporated into synthetic data generation pipelines to create datasets for training and advancing models in various sectors, including healthcare, retail, and manufacturing. These developments demonstrate the potential of LLMs to generate meaningful synthetic data across different domains. Although the integration of LLMs for data augmentation in ML workflows is still in its early stages, their ability to generate synthetic data based on learned patterns from existing datasets offers a promising solution, particularly when data is limited for specific ML tasks Geng et al. (2023).
@@ 3.2.3 Feature Synthesis
LLMs have proven highly effective in generating semantically and contextually relevant synthetic features, enhancing the feature synthesis process through an interpretable, human-in-the-loop approach. In the Context-Aware Automated Feature Engineering (CAAFE) framework Hollmann et al. (2024), LLMs iteratively generate additional meaningful features for tabular datasets based on the dataset description, as shown in Fig. 3. This approach not only produces Python code for creating new features but also provides explanations of the utility and relevance of the generated features. CAAFE represents a step forward in semi-automated data science tasks, emphasizing the importance of context-aware solutions that extend AutoML systems into more interpretative and human-centered workflows.
@@ 4.1.1 Retrieval-based Model Selection
Additionally, Öztürk et al. (2022) introduces a zero-shot AutoDL method that meta-learns a large pre-trained model to select the best deep learning model from candidates, conditioned on dataset meta-features.
@@ 4.2.2 Prediction-based HPO
Prediction-based HPO refers to methods where LLMs recommend hyperparameter settings based on predicted performance, bypassing the need for actual model training and evaluation. This approach allows for faster and more resource-efficient optimization. An example of such a system is AgentHPO Liu et al. (2024a), a two-agent framework consisting of a Creator and Executor agent, which leverages LLM-powered autonomous agents to overcome the complexities of traditional AutoML methods. The Creator agent initiates the optimization process by allowing users to input task-specific details in natural language. It then interprets the input and generates initial hyperparameters, mimicking the expertise of a human specialist. The Executor agent takes the hyperparameters generated by the Creator and handles tasks like training models, recording experimental results, and conducting outcome analysis. The Creator agent further iteratively refines the hyperparameters based on the insights gleaned from the Executor’s training history, streamlining the optimization process and making it more intuitive and efficient. Furthermore, AutoML-GPT Zhang et al. (2023d) predicts model performance by generating a training log for a given hyperparameter configuration based on the provided data card and model card. The system then tunes the hyperparameters by analyzing the predicted training log, which includes relevant metrics and information from the simulated training process.
@@ 6.1.4 Interpretability
To address these challenges, future research should focus on integrating explainable AI (XAI) techniques into LLM-driven workflows. XAI methods Hoffman et al. (2018) could offer deeper insights by complementing textual explanations with visual tools, causal diagrams, or interactive features that highlight how specific features or parameters affect model performance. Furthermore, incorporating human-in-the-loop mechanisms Hollmann et al. (2024) would ensure that users are not merely passive recipients of information but actively engaged in confirming the validity and appropriateness of the system’s decisions.
@@ 6.1.6 Social Impact
However, this automation also presents challenges. One of the key concerns is the potential displacement of human workers. As systems become more autonomous and capable of managing complex ML tasks, there is a risk that they could diminish or replace roles traditionally held by data scientists, engineers, and researchers. This shift could have significant ramifications for employment within the field of ML and beyond Huang et al. (2024). For instance, CAAFE Hollmann et al. (2024) illustrates how automating routine data science tasks could lead to a reduction in the demand for manual labor, potentially displacing workers in data-centric industries. Nevertheless, it also emphasizes that such automation could free up human professionals to focus on more strategic decision-making, enhancing their productivity. Moving forward, developing frameworks where AI augments rather than replaces human labor could mitigate these concerns by encouraging a human-in-the-loop approach to maintain human oversight and creativity Huang et al. (2024).
@@ 6.2 Future Directions
Building on the challenges and opportunities discussed, future research directions offer several exciting possibilities for advancing LLM-driven ML workflows. One promising avenue is the development of end-to-end ML workflow construction using LLMs. This would involve LLMs autonomously managing every stage of the ML pipeline—from data preprocessing to model evaluation—thus creating fully automated workflows. The goal is to further minimize human intervention, enabling seamless construction, optimization, and deployment of ML models. However, realizing this vision requires addressing key challenges such as hallucinations, interpretability issues, and computational costs to ensure that the generated workflows are accurate, reliable, and efficient. Therefore, it is crucial to strike a balance with advanced human-in-the-loop frameworks, where human oversight acts as a safeguard, ensuring that the automated decisions made by LLMs are accountable, fair, and aligned with human expectations, especially in high-stakes applications.

### v2

@@ 1 Introduction
These limitations highlight the need for more advanced solutions, where LLMs demonstrate significant potential by incorporating prior knowledge, facilitating human-AI collaboration, and producing more interpretable outcomes. For instance, in hyperparameter optimization Zhang et al. (2023c); Liu et al. (2024a), LLMs can leverage historical data and domain-specific insights to predict optimal configurations, thereby enhancing model performance and reducing the reliance on exhaustive trial-and-error methods. Furthermore, the sophisticated NLP capabilities of LLMs enable them to act as interactive agents or chatbots, capable of generating and adjusting AutoML code based on contextual information provided by users Arteaga Garcia et al. (2024); Guo et al. (2024); Dakhel et al. (2023). Crucially, in the process of conversational construction of ML workflows—such as feature engineering Hollmann et al. (2024)—LLMs have shown the ability to generate human-readable and explainable features, offering a level of transparency and guidance that surpasses many traditional black-box AutoML systems Zhang et al. (2024a); Nam et al. (2024).
However, integrating LLMs into ML workflows also presents several challenges. Issues such as reasoning hallucinations, ethical concerns, and the substantial computational demands of deploying large-scale models remain significant barriers to their widespread adoption Bommasani et al. (2021); Hollmann et al. (2024); Yao et al. (2024a). Additionally, while LLMs have demonstrated impressive results in specific tasks, their effectiveness across the entire spectrum of ML workflow stages Zhang et al. (2023b); Schmidt et al. (2024) still requires thorough investigation and evaluation using some benchmarks like MLE-Bench Chan et al. (2024) and MLAgentBench Huang et al. (2024).
This survey provides a comprehensive overview of the current state of research on the application of LLMs in constructing and optimizing ML workflows. To the best of our knowledge, it is the first survey to systematically address every stage of the ML workflow (Fig. 1), distinguishing it from previous reviews, such as Tornede et al. (2023a), which primarily explore the broader opportunities of an integration of LLMs and AutoML. We organize recent papers based on the specific steps in which LLMs have been utilized, as summarized in Table 1, with entries sorted by publication year (oldest to newest) and the order of workflow steps. Notably, our focus is on capturing the breadth and possibility of research within the specific context of constructing and optimizing ML workflows, rather than providing an exhaustive examination of all LLM-related methods at each individual stage.
@@ 3.1.1 Data Acquisition
In the context of dataset recommendation, some systems harness LLMs to suggest datasets that align with specific task needs. AutoMMLab Yang et al. (2024) employs LLMs for natural language interaction with users, utilizing a "dataset zoo" to recommend or retrieve relevant datasets based on task descriptions. This automated selection helps tailor custom datasets to the unique requirements of specialized machine learning tasks. Similarly, AutoML-Agent Trirat et al. (2024) employs an LLM-based Data Agent to execute sub-tasks within a decomposed plan generated through retrieval-augmented planning. These sub-tasks include dataset retrieval, pre-processing, augmentation, and analysis, ensuring that the selected datasets align seamlessly with the given task requirements. LLMs are also used for dataset summarization, aiding users in understanding data structure and content before further processing. The Virtual Interactive Data Scientist (VIDS) framework Hassan et al. (2023), for example, features ChatGPT-powered agents that extract detailed insights, including dataset structures, column details, and visualization suggestions. Similarly, JarviX Liu et al. (2023) incorporates an "Insight" component to automatically gather structured data information—such as column names, types, and statistics—and generate LLM-driven summary reports, giving users a quick overview of key dataset characteristics.
@@ 3.1.3 Data Augmentation
Data augmentation refers to techniques used to artificially increase the size of the training dataset by creating enhanced versions of existing data. This step is particularly useful for enhancing model robustness, improving generalization, and avoiding overfitting, especially in scenarios where data is scarce.
LLMs contribute significantly to data augmentation by employing techniques for modifying original data and generating synthetic datasets. In the realm of original data modification, LLMs assist in applying diverse transformation techniques to augment data effectively. For instance, the Data Agent in AutoML-Agent Trirat et al. (2024) suggests various image transformation methods, such as random horizontal flipping and zooming, to enhance user-input image datasets, thereby increasing data variability and robustness for model training. For synthetic data generation, recent studies underscore the capabilities of pre-trained language models (PLMs) to create synthetic datasets for training target models Gao et al. (2023); Yu et al. (2024). To address potential biases in synthetic data generated by a single PLM, FuseGen Zou et al. (2024) innovatively combines multiple PLMs to collaboratively produce higher-quality synthetic datasets without incurring additional queries to individual models, thereby improving the reliability of synthetic data. Similarly, NVIDIA’s Nemotron-4 340B model family Adler et al. (2024) has been integrated into synthetic data generation pipelines to produce datasets for diverse applications, including healthcare, retail, and manufacturing. Although the integration of LLMs into data augmentation is still in its early stages, their ability to modify and generate meaningful data based on learned patterns from existing datasets presents a promising avenue. This is particularly valuable in scenarios where data availability is limited, enabling more effective solutions for specific ML tasks Geng et al. (2023).
@@ 3.2.3 Feature Synthesis
LLMs have proven highly effective in generating semantically and contextually relevant synthetic features, enhancing the feature synthesis process through an interpretable, human-in-the-loop approach. In the Context-Aware Automated Feature Engineering (CAAFE) framework Hollmann et al. (2024), LLMs iteratively generate additional meaningful features for tabular datasets based on the dataset description, as shown in Fig. 3. This approach not only produces Python code for creating new features but also provides explanations of the utility and relevance of the generated features. CAAFE represents a step forward in semi-automated data science tasks, emphasizing the importance of context-aware solutions that extend AutoML systems into more interpretative and human-centered workflows Lindauer et al. (2024).
@@ 4.1.1 Retrieval-based Model Selection
Additionally, the Model Agent within the AutoML-Agent framework Trirat et al. (2024) performs model retrieval and hyperparameter suggestion, guided by insights provided by an Agent Manager regarding high-performing models and suitable hyperparameters for the specific ML task. Moreover, Öztürk et al. (2022) introduces a zero-shot AutoDL method that meta-learns a large pre-trained model to select the best deep learning model from candidates, conditioned on dataset meta-features.
@@ 4.2.2 Prediction-based HPO
Prediction-based HPO refers to methods where LLMs recommend hyperparameter settings based on predicted performance, bypassing the need for actual model training and evaluation Trirat et al. (2024). This approach allows for faster and more resource-efficient optimization. An example of such a system is AgentHPO Liu et al. (2024a), a two-agent framework consisting of a Creator and Executor agent, which leverages LLM-powered autonomous agents to overcome the complexities of traditional AutoML methods. The Creator agent initiates the optimization process by allowing users to input task-specific details in natural language. It then interprets the input and generates initial hyperparameters, mimicking the expertise of a human specialist. The Executor agent takes the hyperparameters generated by the Creator and handles tasks like training models, recording experimental results, and conducting outcome analysis. The Creator agent further iteratively refines the hyperparameters based on the insights gleaned from the Executor’s training history, streamlining the optimization process and making it more intuitive and efficient. Furthermore, AutoML-GPT Zhang et al. (2023d) predicts model performance by generating a training log for a given hyperparameter configuration based on the provided data card and model card. The system then tunes the hyperparameters by analyzing the predicted training log, which includes relevant metrics and information from the simulated training process.
@@ 6.1.4 Interpretability
To address these challenges, future research should focus on integrating explainable AI (XAI) techniques into LLM-driven workflows. XAI methods Hoffman et al. (2018) could offer deeper insights by complementing textual explanations with visual tools, causal diagrams, or interactive features that highlight how specific features or parameters affect model performance. Furthermore, incorporating human-in-the-loop mechanisms Hollmann et al. (2024); Lindauer et al. (2024) would ensure that users are not merely passive recipients of information but actively engaged in confirming the validity and appropriateness of the system’s decisions.
@@ 6.1.6 Social Impact
However, this automation also presents challenges. One of the key concerns is the potential displacement of human workers. As systems become more autonomous and capable of managing complex ML tasks, there is a risk that they could diminish or replace roles traditionally held by data scientists, engineers, and researchers. This shift could have significant implications for employment within the field of ML and beyond Huang et al. (2024). For instance, CAAFE Hollmann et al. (2024) illustrates how automating routine data science tasks could lead to a reduction in the demand for manual labor, potentially displacing workers in data-centric industries. Nevertheless, it also emphasizes that such automation could free up human professionals to focus on more strategic decision-making, enhancing their productivity. Moving forward, developing frameworks where AI augments rather than replaces human labor could mitigate these concerns by encouraging a human-in-the-loop approach to maintain human oversight and creativity Huang et al. (2024).
@@ 6.2 Future Directions
Building on the challenges and opportunities discussed, future research directions present exciting possibilities for advancing LLM-driven ML workflows.
One promising direction is the development of end-to-end ML workflow construction using LLMs, where LLMs autonomously handle every stage of the ML pipeline—from data preprocessing to model evaluation—creating fully automated workflows. The goal is to further reduce human intervention, enabling seamless construction, optimization, and deployment of ML models. Some existing agent-based methods, such as AIDE Schmidt et al. (2024), ResearchAgent Huang et al. (2024), CodeActAgent Wang et al. (2024e, f), and DS-Agent Guo et al. (2024), already offer implicit stage-wise solutions for end-to-end ML workflow generation. These approaches leverage the generative capabilities of LLMs to produce solution drafts and iteratively refine them based on performance feedback and interactions with the environment. However, this process often lacks transparency, making it difficult for users to personalize or modify individual workflow components. In contrast, explicit end-to-end ML workflow generation could provide a complete ML solution while allowing users to control and customize each component.